OpenStudy (anonymous):

45*-45*-90* Right Triangles Help please? Find the hypotenuse of each isosceles right triangle when the legs are of the given measure. Given = 3 sqrt 2 Inches = ?

jhonyy9 (jhonyy9):

so in case of a right triangle for you can calcule a side just you need using pythagora do you know it sure

jhonyy9 (jhonyy9):

so isoscele triangle mean that two sides are equales yes ?

jhonyy9 (jhonyy9):

hypothenuse ^2 =a^2 +b^2 but because a=b so than h^2 =2a^2 and because a=b=3sqrt2 so than h^2 =2*(3sqrt2)^2 h^2 =2*9*2 h^2 =36 h =sqrt36 =6 yes is true that h=+/- 6 but you know that a lengh of a side not can being minus value h=6

jhonyy9 (jhonyy9):

hope that is understandably good luck bye

OpenStudy (anonymous):

oh sorry i was taking a test crap i feel all rude now lol thank you soo much!

jhonyy9 (jhonyy9):

yw good luck bye no problem